### v2

# Initial-State Typicality in Quantum Relaxation

Abstract
Relaxation in open quantum systems is fundamental to quantum science and technologies. Yet, the influence of the initial state on relaxation remains a central, largely unanswered question. Here, by systematically characterizing the relaxation behavior of generic initial states, we uncover a typicality phenomenon in high-dimensional open quantum systems: relaxation becomes nearly initial-state-independent as system size increases under verifiable conditions. Crucially, we prove this typicality for many thermalization processes above a size-independent temperature. Our findings extend the typicality to transient open quantum dynamics, in turn identifying a class of systems where two widely used quantities—the Liouvillian gap and the maximal relaxation time—merit re-examination. We formalize this with two new concepts: the “typical strong Mpemba effect” and the “typical relaxation time”. Beyond these conceptual advances, our results provide practical implications: a scalable route to accelerating relaxation and a typical mixing-time benchmark that complements conventional worst-case metrics for quantum simulations and state preparation.
Introduction—Relaxation is a ubiquitous phenomenon in nature and a central process in open quantum systems. The relaxation time, also known as the mixing time, is a key figure of merit in diverse areas, including quantum information and computing [1, 2, 3], quantum simulations [4, 5, 6, 7], state preparation [8, 9, 10, 11], and quantum heat engines [12, 13, 14, 15, 16]. Over the past decades, influential lines of work employing rigorous mixing-time [17, 18, 19, 20, 21, 22, 23, 24, 25, 26] and spectral analysis [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39] have clarified how the structure of the dynamics controls relaxation in open systems. These approaches have been instrumental in characterizing remarkable phenomena such as the cutoff phenomenon in classical and quantum Markov processes [17, 18], the Liouvillian skin effect [32, 33] and the rapid mixing [19, 20, 21, 22, 23, 24, 25]. A unifying aspect of these studies is their focus on identifying a single, initial-state-independent timescale to characterize a given dynamics. However, this approach overlooks the conceptually and practically important influence of the initial state on relaxation. Since every relaxation process begins from a specific state, its impact cannot be disregarded.
A shift of perspective emerged with the Markovian Mpemba effect established in [27, 28]: under fixed dynamics, different initial states can relax at dramatically different rates. The Mpemba effect has been intensively studied in classical systems [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50], and inspired other studies on anomalous initial-state effect [51, 52, 53]. Follow-up works extended the Mpemba effect to quantum systems [29, 54, 30, 55, 56, 57, 58, 31, 59, 60, 61, 62, 63, 64, 65, 66, 67], inaugurating a line of initial-state engineering for accelerating relaxation. See [68, 69] for recent reviews. Despite these advances, a key question remains open: moving beyond the specific, pre-selected initial states and small systems common in Mpemba-effect studies, is there a universal relaxation behavior for generic initial states, how does it scale with system size, and—if affirmative—what scalable practical utility does it enable?
Here we address this question by proposing a theoretical framework for characterizing the universal relaxation behavior of generic initial states across different system sizes. Crucially, we uncover a typicality phenomenon (concentration) in the relaxation of high-dimensional open quantum systems. Typicality is of fundamental importance and is regarded as a cornerstone of statistical mechanics [70, 71]. To date, it has been established mainly for equilibrium quantum states [70, 71, 72], closed quantum dynamics [73, 74, 75, 76], quantum machine learning [77, 78] and classical wave systems [79], while its counterpart for transient open quantum dynamics has remained largely unknown. We show that, under a large class of conditions, the relaxation dynamics becomes insensitive to the initial state. We provide explicit examples where these conditions are satisfied, including spin chains, rapid-mixing systems and high-temperature thermalization. Insights from random matrix theory suggest the potential universality of this phenomenon. Moreover, the phenomenon prompts a re-evaluation of two well-established concepts: the Liouvillian gap and the maximal relaxation time.
In addition to these new fundamental insights on open quantum dynamics, our results yield practical utility: the typicality phenomenon implies that simple randomized preparations via techniques like the unitary 2-design [80] may suffice to achieve fast relaxation. This offers a scalable route to accelerating quantum relaxation that avoids fine-tuning and detailed prior knowledge of the dynamics, thereby directly relevant to large-scale quantum platforms. Furthermore, our work provides a novel typical-case efficiency benchmark for dissipative quantum tasks, such as the evaluation of quantum Gibbs samplers [7].
Setup—We consider a Markovian open quantum system living in a Hilbert space $\mathcal{H}$ of dimension $d$, whose dynamics is described by the Lindblad-Gorini-Kossakowski-Sudarshan master equation
$\frac{d\rho}{dt}=\mathcal{L}(\rho)\equiv-i[H,\rho]+\sum_{i}\left[J_{i}\rho J_{i}^{\dagger}-\frac{1}{2}\{J_{i}^{\dagger}J_{i},\rho\}\right].$ (1)
The superoperator $\mathcal{L}$ is referred to as Lindbladian throughout this work. In the above, $H$ is the Hamiltonian of the system and the jump operators $J_{i}$ describe the dissipative coupling to the environment, [A,B]:=ABBA[A,B]:=AB-BA is the commutator and $\{A,B\}:=AB+BA$ denotes the anticommutator for any operators A,BA,\ B. The eigenvalues of $\mathcal{L}$ are denoted $\lambda_{k}$ and ordered such that $0=\lambda_{1}\geq\text{Re}(\lambda_{k})\geq\text{Re}(\lambda_{k+1}),\ k\geq 2$. Assuming that $\mathcal{L}$ is diagonalizable, we define $R_{k}$ and $L_{k}$ as the right and left eigenmatrices corresponding to $\lambda_{k}$, satisfying $\mathcal{L}(R_{k})=\lambda_{k}R_{k}$ and $\mathcal{L^{\dagger}}(L_{k})=\lambda_{k}^{*}L_{k},\ k=1,...,d^{2}$, with the dual superoperator $\mathcal{L}^{\dagger}$ governing the Heisenberg-picture dynamics of observables:
$\mathcal{L}^{\dagger}(O)=i[H,O]+\sum_{i}\left[J_{i}^{\dagger}OJ_{i}-\frac{1}{2}\{J_{i}^{\dagger}J_{i},O\}\right].$
We normalize the eigenmatrices by the biorthogonal condition
${\rm tr}(L_{k}^{\dagger}R_{h})\equiv\delta_{kh},\quad\|R_{k}\|_{1}=1,$ (2)
where $\|A\|_{1}:={\rm tr}(\sqrt{A^{\dagger}A})$ is the trace norm. The additional constraint $\|R_{k}\|_{1}=1$ excludes any size-dependence from $R_{k}$ on the relaxation. We assume $\mathcal{L}$ has a unique stationary state $\rho_{\rm ss}$ (however, our analysis can be directly extended to cases where multi-stability exists). Then, given an arbitrary initial state $\rho_{0}$, the system density matrix at time $t$ is
$\rho(t)=e^{t\mathcal{L}}(\rho_{0})=\rho_{\rm ss}+\sum_{k=2}^{d^{2}}e^{\lambda_{k}t}a_{k}R_{k},$ (3)
where coefficients $a_{k}\equiv{\rm tr}(L_{k}^{\dagger}\rho_{0})$ determine the overlaps between the $k$th relaxation mode and the initial state.
In the long-time limit, the density matrix evolves as
$\|\rho(t)-\rho_{\rm ss}\|_{1}\overset{\text{large }t}{\sim}\begin{cases}|e^{\lambda_{2}t}a_{2}|,&\text{real $\lambda_{2}$},\\[6.0pt] |e^{\lambda_{2}t}a_{2}|+|e^{\lambda_{2}^{*}t}a_{2}^{*}|,&\text{others},\end{cases}$ (4)
where we use $\|R_{2}\|_{1}=\|R_{2}^{\dagger}\|_{1}=1$. From Eq. (4), it is clear that the relaxation timescale is determined by the slowest mode, i.e., $\tau_{\rm rel}\sim 1/|{\rm Re}(\lambda_{2})|$. Eq. (4) also implies that, the relaxation starting from an initial state with a smaller $|a_{2}|$ will be faster than that from one with a larger $|a_{2}|$  [^1] This can be more clearly seen by defining an $\epsilon$-relaxation time $\tau^{\epsilon}_{\rm rel}$ as in [32]. It is defined by the relation $|a_{2}|e^{-|{\rm Re}(\lambda_{2})|\tau^{\epsilon}_{\rm rel}}=\epsilon$, which gives $\tau^{\epsilon}_{\rm rel}=\frac{1}{|{\rm Re}(\lambda_{2})|}\log(|a_{2}|/\epsilon)$.. In particular, if $a_{2}=0$, the relaxation timescale becomes $1/|{\rm Re}(\lambda_{3})|$, termed the strong Mpemba effect (SME) [28]. While the slowest mode, which has been the sole focus of most previous studies, governs the long-time dynamics, at intermediate times other modes $(k>2)$ can also contribute, motivating a more complete view.
The existing literature has largely focused on specific, engineered initial states, such as those tailored to produce the SME [29, 30]. This leaves a more fundamental question unanswered: What is the relaxation behavior for a generic initial state chosen from the Hilbert space? Answering this requires a shift from analyzing specific values of $a_{k}$ to understanding their collective statistical properties.
Typical relaxation behaviors with respect to the initial state—We now address this question under a fixed $\mathcal{L}$. To make the notion of a “generic state” concrete, we model it by considering states $\rho_{0}$ drawn from a few representative random ensembles and analyzing the resulting statistics of their coefficients $a_{k}(\rho_{0})$. In this setting, the overlaps $a_{k}$ become random variables, while $L_{k},\ R_{k}$ and $\lambda_{k}$ remain fixed, allowing us to characterize the typical initial-state dependence of relaxation. For each ensemble, we focus on the first two moments of this distribution due to their experimental feasibility.
We begin by examining an ensemble generated by an operationally convenient procedure. We start with any convenient initial state $\rho^{\prime}_{0}$ and apply a random unitary (from a unitary $2$-design) to prepare the initial state for the relaxation, denoted as $\rho_{0}=U\rho_{0}^{\prime}U^{\dagger}$. When $\rho^{\prime}_{0}$ is a pure state, the situation reduces to having the initial state of the relaxation process of interest randomly sampled from a distribution that precisely matches the first two moments of the Haar measure. The variance of $a_{k}={\rm tr}(L_{k}^{\dagger}\rho_{0})$ over the random unitary $U$ in a 2-design is given by (see Supplemental Material [82] for the derivation)
${\rm Var}_{U}(a_{k})=\frac{[{\rm tr}(\rho_{0}^{\prime 2})-\frac{1}{d}]\left[\|L_{k}\|_{2}^{2}-\frac{|{\rm tr}(L_{k}^{\dagger})|^{2}}{d}\right]}{d^{2}-1},$ (5)
where the variance is defined as ${\rm Var}(a_{k}):=\langle|a_{k}|^{2}\rangle-|\langle a_{k}\rangle|^{2}$ and $\|A\|_{2}:=\sqrt{{\rm tr}(A^{\dagger}A)}$ is the Hilbert-Schmidt (HS) norm. The inequalities $|{\rm tr}(L_{k}^{\dagger})|^{2}\leq\|\mathbb{I}_{d}\|^{2}_{2}\|L_{k}\|_{2}^{2}=d\|L_{k}\|_{2}^{2}$ and ${\rm Tr}(\rho_{0}^{2})\geq 1/d$ guarantee that the variance is non-negative. To proceed, we introduce the diagonal eigenvalue condition number [83, 84] $O_{k}:=\|L_{k}\|_{2}\|R_{k}\|_{2}$, which controls the sensitivity of $\lambda_{k}$ to perturbations [83]. The HS norm and the trace norm are connected via the relation $\|A\|_{2}\leq\|A\|_{1}\leq\sqrt{d}\|A\|_{2}$. Consequently, the variance is upper bounded as ($\|R_{k}\|_{1}=1$)
${\rm Var}_{U}(a_{k})<\frac{\|L_{k}\|_{2}^{2}}{d^{2}}\leq\frac{O_{k}^{2}}{d}.$ (6)
Thus, if $O_{k}$ grows slower than $\sqrt{d}$, the variance of $a_{k}$ converges to 0 in the $d\to\infty$ limit, e.g. $O_{k}\sim d^{a},\ a\in[0,1/2)$ or grows logarithmatically. That is, the variance exponentially decays to zero as the system size increases (e.g., for a one-dimensional quantum chain consisting of $n$ qubits, $d=2^{n}$) and $a_{k}$ concentrates sharply around its mean value,
$\langle a_{k}\rangle=\frac{{\rm tr}(L_{k}^{\dagger})}{d}.$ (7)
That is, most initial states have the same $a_{k}\approx\langle a_{k}\rangle$, though they are not necessarily close to each other. This establishes a sufficient condition of concentration of the $k$th relaxation mode.
Notably, $O_{k}$ has a clear physical interpretation: it measures the robustness of the $k$th mode’s characteristic timescale against perturbations on the Lindbladian. A large $O_{k}$ signifies that the mode is physically fragile, as its timescale $1/|{\rm Re}(\lambda_{k})|$ is highly sensitive to small uncertainties or perturbations within the system’s Hamiltonian or dissipative channels, see  [^2] For a perturbation $\delta\mathcal{L}$ to the Lindbladian, the resulting first-order change in the eigenvalue $\lambda_{k}$ is given by $\delta\lambda_{k}=\text{Tr}(L_{k}^{\dagger}\delta\mathcal{L}R_{k})$. This leads to the bound $|\delta\lambda_{k}|\leq O_{k}\|\delta\mathcal{L}\|_{\rm op}$, where $\|\cdot\|_{\rm op}$ is the operator norm. See [82, 83] for more details. Thus, relaxation modes that are more robust to perturbations are easier to concentrate.
The random mixed state generated by performing a random unitary on a mixed state, may be appealing for further experimental verifications and applications due to its relative operational simplicity. However, such mixed states are not fully random, because unitary operations do not affect the eigenvalues of a density matrix, so their spectrum is not random. We thus need to consider random mixed states induced by the partial trace to ensure true randomness. For an initial state $\rho_{0}$ randomly selected from a truly random mixed state ensemble (the HS measure [86]), the variance scales as [82]
${\rm Var}(a_{k})=\frac{\|L_{k}\|_{2}^{2}-\frac{|{\rm tr}(L_{k}^{\dagger})|^{2}}{d}}{d(d^{2}+1)}<\frac{\|L_{k}\|_{2}^{2}}{d^{3}}\leq\frac{O_{k}^{2}}{d^{2}},$ (8)
while the average $\langle a_{k}\rangle$ is still given by Eq. (7). Thus, for a random mixed state, the sufficient condition for concentration weakens to $O_{k}$ grows slower than $d$.
A more meticulous description of the concentration phenomenon is the probability of deviating from mean value. Using Chebyshev’s inequality, we obtain
$P(|a_{k}-\langle a_{k}\rangle|>\epsilon)\leq\frac{{\rm Var}(a_{k})}{\epsilon^{2}}<\frac{\|L_{k}\|_{2}^{2}}{d^{q}\epsilon^{2}}\leq\frac{O_{k}^{2}}{d^{q-1}\epsilon^{2}},$ (9)
where $q=2$ for random pure states and $q=3$ for random mixed states. This offers a more direct intuition of concentration: in the thermodynamic limit, most initial states yield a value of $a_{k}$ that lives within a small region around the mean.
On the other hand, if $\|L_{k}\|_{2}$ grows faster than $\mathcal{O}(d^{3/2})$, the variance will diverge in the $d\to\infty$ limit for both pure and mixed random initial states. There are known examples where this is the case: the Liouvillian skin effect [32] and the super-exponential overlap coefficients [34]. This drastic change may be termed the “typicality transition”.
So far, we require a “generic” initial state not to have any preference (uniformly random), thereby faithfully reflecting the underlying properties of $\mathcal{L}$. However, we demonstrate in [82] that the concentration can still occur with an arbitrary physical constraint (or, prior knowledge) on the initial state.
The scaling conditions for the appearance of concentration of quantum relaxation modes or divergence of their variances are a main prediction of this work. If the concentration conditions are satisfied, the initial state becomes irrelevant for relaxation dynamics as the dimension $d$ increases. To be more specific, when $a_{2}$ concentrates, the long-time relaxation dynamics concentrates, and the relaxation time becomes irrelevant to the initial state; if other $a_{k}$ also concentrate, the full relaxation dynamics concentrates and even the early and intermediate-time dynamics are nearly independent of the initial state. This novel concentration phenomenon in the transient dynamics of open quantum systems is a complementary finding to results on the typicality in equilibrium states [70, 72, 71] and closed quantum dynamics [75, 76].
To provide a concrete understanding of the concentration phenomenon, we now present explicit examples demonstrating how the dimensional scaling of $O_{k}$ or $\|L_{k}\|_{2}$ meets the established requirement for concentration. We defer the discussion of some important implications and practical applications to the end of this work.
Example 1: Non-interacting dissipative $N$-qubit chain—We consider a non-interacting system consisting of $N$ damping qubits, with Hamiltonian $H=\sum_{i=1}^{N}E\sigma^{z}_{i}/2$ and local jump operators $J_{i,\downarrow}=\sqrt{\gamma_{1}}\,\sigma_{i}^{-},\quad J_{i,\uparrow}=\sqrt{\gamma_{0}}\;\sigma_{i}^{+}$. Here, $\sigma^{z}_{i}$ are Pauli operators of the $i$th particle, $\sigma_{i}^{-}=(|0\rangle\langle 1|)_{i}$ and $\sigma_{i}^{+}=(|1\rangle\langle 0|)_{i}$. The system can be analytically solved to illustrate the concentration phenomenon. In this case, it can be verified that [82] $\|L_{k}\|_{2}\sim\mathcal{O}(2^{N/2})\sim\mathcal{O}(\sqrt{d})$. Thus, the variances for both ensembles exponentially converge to 0 as ${\rm Var}_{U}(a_{2})\sim d^{-1}\sim 2^{-N},\ {\rm Var}(a_{2})\sim d^{-2}\sim 4^{-N}$. One can further add the dephasing noise $\sigma_{i}^{z}$, which will not change the scaling.
Example 2: Dissipative transverse-field Ising model (TFIM)—We then numerically study an open TFIM of length $N$ with local dissipation. The Hamiltonian is $H=-J\sum_{i=1}^{N-1}\sigma_{i}^{z}\sigma_{i+1}^{z}-g\sum_{i=1}^{N}\sigma_{i}^{x}$, and jump operators are the same as Example 1 (with a physical constraint $\gamma_{1}/\gamma_{0}=e^{\beta E},$ where $\beta=1/k_{B}T$ is the inverse temperature of multiple identical heat baths independently coupled to each qubit). Notably, $\rho_{\rm ss}$ may not be an equilibrium state in this example. We find that the concentration of $a_{2}$ is robust in this model, across different parameter regimes ($J,\ g,\ \beta$), the details of a specific case is shown in End Matter. The robustness is reminiscent of the fact that there is no thermal phase transition in one-dimensional spin systems and there may be a connection to be explored.
Sufficiently rapid relaxation may imply mode concentration—Rapid mixing, a well-known phenomenon states that the relaxation timescale for a system with size $n$ is of the order $\mathcal{O}(\log[{\rm poly}(n)])$, can imply the concentration of $a_{2}$. We prove that $\|L_{k}\|_{2}\leq 2\sqrt{d}\max_{\rho_{0}}|a_{k}|$ for any $k$. This inequality yields upper bounds for the variance: ${\rm Var}_{U}(a_{k})\leq 4(\max_{\rho_{0}}|a_{k}|)^{2}d^{-1}$ and ${\rm Var}(a_{k})\leq 4(\max_{\rho_{0}}|a_{k}|)^{2}d^{-2}$. Indeed, rapid mixing implies $\max_{\rho_{0}}|a_{2}|$ is at most of the order $\mathcal{O}({\rm poly}(\log d))$ under an assumption used in [34], ensuring the variances vanish as $d\to\infty$. The rapid mixing has been proven for diverse systems [19, 20, 21, 22, 24, 87]. Interestingly, rapid mixing implies stability of relaxation [19], and concentration also implies more stable relaxation modes, which indicates deep connections between them.
However, fast relaxation is not a necessary condition for typicality. Since the closing of Liouvillian gap is sufficient for a slow relaxation, the slow relaxation alone cannot imply $O_{k}$ grows fast. Notably, a recent work provided an example for the typicality in slow relaxation [88] with a different normalization $\|R_{k}\|_{2}=1$. In their case, $\|L_{2}\|_{2}\approx\|\mathbb{I}_{d}\|=\sqrt{d}$ and thus ${\rm Var}(a_{2})\sim d^{-1}$.
High-temperature thermalization is sufficient for initial-state typicality—Importantly, we prove that for a large class of thermalization processes with a sufficiently high environment temperature, all relaxation modes concentrate. To be specific, for the Lindbladian (1) obeying the quantum detailed balance condition (QDBC) with respect to a full-rank Gibbs state $\rho_{\beta}:=e^{-\beta H_{0}}/{\rm tr}(e^{-\beta H_{0}})$ with a Hamiltonian $H_{0}$ and an inverse temperature $\beta\equiv 1/(k_{B}T)$, we prove an upper bound for $O_{k}$ corresponding to any $\lambda_{k}$ [82]:
$O_{k}\leq\sqrt{\frac{p_{\rm max}(\rho_{\beta})}{p_{\rm min}(\rho_{\beta})}}\equiv e^{\beta\Delta E/2},$ (10)
where $p_{\max/\min}(\rho_{\beta})$ are maximum/minimum eigenvalues of $\rho_{\beta}$ and $\Delta E:=E_{\rm max}-E_{\rm min}$ is the energy range of $H_{0}$. Here, by QDBC we refer to the Davies map [89, 90] or the Kubo–Martin–Schwinger (KMS) condition [22] with respect to $\rho_{\beta}$  [^3] See [82] for details of these conditions. It has been shown that, if a Lindbladian satisfies these conditions, $\rho_{\beta}=e^{-\beta H_{0}}/{\rm tr}(e^{-\beta H_{0}})$ is its steady state. Note that $H_{0}$ may not coincide with the system Hamiltonian $H$ of the Lindbladian, there could be a shift term, e.g., the Lamb shift [24].. Lindbladians satisfying the KMS condition have been constructed [8, 92] and prove critical to quantum simulations [7].
For most Hamiltonians without long-range interactions, $\Delta E\sim\mathcal{O}(n)\equiv cn$, where $c$ is a size-independent constant and $n$ is the system size. This requirement reflects the extensivity condition, which can also be fulfilled by long-range Hamiltonians when the Kac prescription is applied [93]. Then, take the many-qubit system as an example, if the temperature satisfies the condition $e^{\beta c}\leq 2\Rightarrow\beta<\beta_{c}:=\log 2/c$, i.e., higher than a threshold irrelevant to the system size, $O_{k}$ has a smaller scaling than $\mathcal{O}(\sqrt{d})$ and all $a_{k}$ will concentrate around their mean for an initial Haar ensemble. For the HS ensemble, the threshold turns to $\beta_{c}^{\prime}=2\log 2/c$.
Universality of the typicality for random mixed states—Beyond specific examples, we here provide general arguments, using insight from random matrix theory. For a large class of non-Hermitian random matrix such as complex Ginibre ensemble and random matrix whose entries are independent and identically distributed, the typical scaling of $O_{k}$ for bulk eigenvalues and large $d$ was rigorously established as $O_{k}\sim\mathcal{O}(d)$ [94, 95]. For random Lindbladian, this scaling has also been confirmed [96, 97]. This cannot guarantee the universality of the typicality for pure initial states.
However, if the initial state is a random mixed state, we expect that at least the concentration of the dominant relaxation mode ($a_{2}$) would be a universal behavior. The reason is that $\lambda_{2}$ should be classified as the edge eigenvalue. Indeed, for many random matrix models, the typical scaling of $O_{k}$ of an edge eigenvalue $\lambda_{k}$ was shown to be $\mathcal{O}(\sqrt{d})$ [98, 99]. Although for random Lindbladian, currently there is no rigorous proof, a typical scaling slower than $d$ can still be anticipated, which at least assures the concentration of $a_{2}$ for random mixed states. It is known that edge eigenvalues have larger eigenvalue spacing, which makes them more stable (smaller condition number). In a recent paper [100], it was shown that a larger eigenvalue spacing implies a smaller dimensional scaling of the condition number $O_{k}$.
Two important implications of the typicality—First, the Liouvillian gap $1/|{\rm Re}(\lambda_{2})|$ may not correctly quantify the relaxation time for a class of systems satisfying two conditions: (i) $a_{2}$ concentrates (ii) $|\langle a_{2}\rangle|=|{\rm tr}(L_{2}^{\dagger})|/d$, vanishes. The second condition is met, for example, if ${\rm tr}(L_{2})=0$ [e.g., for a complex $\lambda_{2}$ of a Davies map [30]]. When both conditions hold, $a_{2}$ concentrates sharply around 0. Our Example 1 also falls into this class. We term this the “typical SME” (TSME), as it implies the SME ($a_{2}\approx 0$) is typical for generic initial states. In this regime, the relevant relaxation timescale may be set by $1/|{\rm Re}(\lambda_{k})|$ with $k\geq 3$, instead of $1/|{\rm Re}(\lambda_{2})|$. We discuss the conditions for this phenomenon in detail in [82]. The TSME should be distinguished from the conventional SME, as it represents the intrinsic property of the underlying dynamics (Lindbladian). In contrast, the SME only reflects the property of a specific initial state.
Second, for systems with initial-state typicality, the relevance of the commonly adopted initial-state-independent maximal relaxation time (maximized over all initial states) is called into question, since it may deviate greatly from the typical relaxation timescale and thus may not be physically representative. In our Example 1, all $a_{k}$ concentrate sharply around their mean, and $\max_{k}\langle a_{k}\rangle\sim\mathcal{O}(1)$. By contrast, the maximal relaxation time [34, 35] is calculated from $\max_{k,\rho_{0}}|a_{k}(\rho_{0})|$, which can scale as $\mathcal{O}(d)$.
Definition: typical relaxation time—We therefore propose the “typical relaxation time”, the relaxation time from representative initializations whose $a_{2}$ concentrate around $\langle a_{2}\rangle$. As an illustrative definition analogous to that used for the maximal relaxation time [32, 34, 35], one can define a typical timescale $\tau^{\epsilon}_{\rm rel}$ via the mean overlap of the slowest mode:
$|\langle a_{2}\rangle|e^{-|{\rm Re}(\lambda_{2})|\tau^{\epsilon}_{\rm rel}}=\epsilon.$ (11)
An alternative and more rigorous definition using the trace distance (the “typical mixing time”) is given in the End Matter. We posit that this typical relaxation time is a more physically relevant quantity in the typicality regime. Notably, it still characterizes the intrinsic properties of the generator $\mathcal{L}$, but it does so by reflecting the statistically dominant behavior rather than an exponentially rare, worst-case scenario.
Potential practical utility: scalable acceleration of relaxation—A key practical utility of initial-state typicality is enabling scalable relaxation acceleration. If a system is initially in a slow-relaxing state, applying a randomization protocol (such as a unitary 2-design) can kick it into the fast-relaxing typical set. The presence of a TSME, ($\langle a_{2}\rangle\to 0$ and ${\rm Var}(a_{2})\to 0$ so that $a_{2}\approx 0$ for most states), is a powerful diagnostic indicating that this typical set is indeed characterized by rapid relaxation.
This randomization can be realized via a unitary 2-design, which only requires modest polynomial resources [77]. This can be applied directly to the system of interest, or used to generate a random mixed state by combining an auxiliary system with the system, performing the random unitary on the composite system, and then tracing out the auxiliary part. The technique has been performed experimentally in quantum many-body systems [101], and its key component, the Clifford circuit, has been applied in systems up to 40 qubits [102]. For our purpose, an approximate $2$-design, implementable via scalable random circuits [103, 104, 105], is sufficient [82]. Remarkably, recent advances indicate that achieving an approximate $2$-design on $n$ qubits via random circuits can be done in circuit depth $\mathcal{O}(\log\log n)$ with ancillas [106, 107, 108], and in $\mathcal{O}(\log n)$ without ancillas [109]. Additionally, suitable hybrid quantum–classical approaches may drive atypical states into the typical-relaxation regime in constant parallel depth when $\mathcal{O}(n)$ ancillas are available  [^4] Concretely, when $\mathcal{O}(n)$ ancilla qubits and intermediate measurements are available, the depolarizing action $\rho_{0}\mapsto(1-p)\rho_{0}+p\,\mathbb{I}/d$ can be realized in constant parallel depth per circuit via a hybrid sampling scheme, demonstrated on superconducting hardware up to 27 qubits [117]. Setting $p=1$ yields $a_{2}(\rho_{0})=a_{2}(\mathbb{I}/d)=\langle a_{2}\rangle$ exactly. Our typicality result guarantees that this value is representative of almost all initial states; thus the unitary 2-design route and this shortcut are equivalent in the figure of merit $a_{2}$. Moreover, initial-state typicality provides an additional advantage in this case: since $\mathbb{I}/d$ lies at the center of the typical set, small imperfections in its preparation still correspond to states within that set, making the protocol intrinsically robust..
Our findings also imply a challenge for protocols that rely on finding atypical states. If $a_{2}$ concentrates to a large $\langle a_{2}\rangle$, then using random searching methods [30] to find a rare $a_{2}\approx 0$ state may be inefficient, as the chance to escape from the typical set is exponentially small for many-body systems [77] . Thus, the search may need to be constrained in the weak-entanglement regime.
Re-evaluation of the efficiency of quantum Gibbs samplers—The mixing time is essential in evaluating the efficiency of a quantum Gibbs sampler [8, 92, 7] for efficient state preparations and quantum simulations. In addition to the conventional (worst-case) metric, it is natural to also report the typical relaxation benchmark—namely the typical relaxation time defined above. With initial-state typicality, establishing efficiency may amount to showing typical rapid mixing, i.e., that the typical relaxation time scales favorably, which is a weaker and more representative requirement than conventional rapid mixing. Even if a sampler is initialized atypically, a light randomization can push it into the typical set, as discussed above.
Conclusion and Outlook—We established the phenomenon of initial-state typicality in the relaxation of open quantum systems. This finding has profound conceptual and practical implications. Conceptually, it challenges the universality of the Liouvillian gap and the maximal relaxation time as faithful descriptors of relaxation. To address this, we introduced the TSME as a diagnostic and proposed the typical relaxation time as a more relevant timescale in the typicality regime. On a practical level, the phenomenon underpins a scalable protocol for relaxation acceleration and offers new ways to assess the efficiency of algorithms like quantum Gibbs samplers.
Our work opens several promising avenues for future research. A concrete next step is to construct explicit physical models that exhibit a sharp transition into the typicality regime. Further questions concern the fluctuation-dominated regime where typicality fails: are there other universal statistical laws governing relaxation, perhaps revealed in the distribution of relaxation times? The potential universality of the typicality itself, hinted at by random matrix theory, also warrants deeper investigation. Additionally, the concentration of slow modes implies that associated metastable states [111] also exhibit typicality, potentially impacting quantum computation that leverages metastability [112, 113]. Perhaps most importantly, the typical relaxation time provides a powerful new tool. It may enable proofs of rapid mixing under broader conditions and redefine performance benchmarks across quantum science, from state preparation to quantum simulations.
Acknowledgments—I am grateful to Zongping Gong, Guohua Xu and Hongchao Li for useful discussions, and Yue Liu for a critical reading of the manuscript. I also acknowledge László Erdös for the kind help in private communications. R. B. is supported by JSPS KAKENHI Grant No. 25KJ0766.